40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 7 18:18:55 2009, Jenne, Update, COC, ETMX drag wiped 
    Reply  Tue Dec 8 13:03:26 2009, Koji, Update, COC, ETMY drag wiped 
Message ID: 2361     Entry time: Mon Dec 7 18:18:55 2009     Reply to this: 2366
Author: Jenne 
Type: Update 
Category: COC 
Subject: ETMX drag wiped 

[Koji, Jenne, Alberto, Steve, Bob]

ETMX has been drag wiped. 

Around 2:45pm, after the main IFO volume had come up to atmospheric pressure, we removed both doors to the ETMX chamber.  Regular procedures (wiping of O-rings with a dry, lint-free cloth, covering them with the light O-ring covers, etc.) were followed.  Koji took several photos of the optic, and the rest of the ETMX chamber before anything was touched. These will be posted to the 40m Picasa page.  Steve and Koji then deionized the optic.

Koji removed the bottom front earthquake stop, and clamped the optic with the remaining earthquake stops.

The clean syringes were prepared: These are all glass and metal (nothing else) medical syringes.  The size used was 100microliters.  Earlier today, we had prepared our solvents in small little beakers which had been baked over the weekend.  Brand new glass bottles of Acetone and Isopropyl Alcohol were opened, and poured into the small beakers.  To make sure we have enough, we have 3 ~10ml beakers of each Acetone and Isopropyl.

We started with Acetone.  The syringe was filled completely with acetone, then squirted onto a kimwipe.  This was repeated ~twice, to ensure the syringe was well rinsed.  Then the syringe was filled a little past the 100 microliter mark.  Koji held a piece of lens cleaning paper to ETMX and used an allen wrench underneath the optic to help guide the paper, and keep it near the optic (of course, the only thing in actual contact with the optic was the lens paper).  In one smooth shot, the plunger of the syringe was pressed all the way down.   (This is a bit tricky, especially when the syringe is totally full.  You have to squeeze it so the plunger moves fairly quickly down the barrel of the syringe to get a good arc of liquid.  The goal is to shoot all of the solvent to the same place on the lens paper, so that it makes a little circle of wetness on the paper which covers the coated part of the optic.  The amount of solvent used should be balanced between having too little, so that the paper is dry by the time it has been wiped all the way down, and too much such that there is still a residue of liquid on the optic after the paper has been removed.)  The target was to hit the optic just above the center mark (the oplev was on, so I went for just above the red oplev dot).  Immediately after applying the liquid onto the paper, Koji slowly and smoothly pulled down on the lens paper until it came off of the bottom of the optic.  The acetone was repeated, for a total of 2 acetone wipes.  Because acetone evaporates very quickly, more acetone is used than isopropyl.  The optimal amount turned out to be ~115 microliters of acetone.  It is hard to say exactly how much I had on the second wipe, because the syringe is not marked past 100 microliters.  On the first wipe, with about 105 microliters, the lens paper was too dry at the bottom of the optic.

We then switched to Isopropyl.  A new syringe was used, and again we rinsed it by filling it completely with isopropyl, and emptying it onto a kimwipe.  This was repeated at least twice.  We followed the same procedure for applying liquid to the optic and wiping the optic with the lens paper.  On the first try with isopropyl, we used 100 microliters, since that was the preferred amount for acetone.  Since isopropyl evaporates much slower than acetone, this was determined to be too much liquid.  On the second isopropyl wipe, I filled the syringe to 50 microliters, which was just about perfect.  The isopropyl wiping was done a total of 2 times.

After wiping, we replaced the front bottom earthquake stop, and released the optic from the other earthquake stops' clamping.  The OSEM values were checked against the values from the screenshots taken yesterday afternoon, and were found to be consistent.  Koji took more photos, all of which will be placed on the 40m Picasa page.

We visually inspected the optic, and we couldn't see anything on the optical surface of the mirror.  Koji said that he saw a few particulates on some horizontal surfaces in the chamber.  Since the optic seemed (at least to the level of human vision without a strong, focused light) to be free of particulates on the optical surface to start with, the suspense will have to remain until we button down, pump down, and try to lock the IFO to determine our new finesse, to see if the wiping helped any substantial amount. 

We replaced the regular, heavy door on the inner side of the ETMX chamber (the side closer to the CES building), and put only a light door on the outer side of the chamber (the side closer to the regular walkway down the arm).  We will look at the spectra of the OSEMS tomorrow, to confirm that none of the magnets are stuck.

We commence at ~9am tomorrow with ETMY.

LESSONS LEARNED:

The LED lights are awesome.  It's easy to use several lights to get lots of brightness (more than we've had in the past), and the chamber doesn't get hot.

We should get larger syringes for the acetone for the large optics.  It's challenging to smoothly operate the plunger of the syringe while it's so far out.  We should get 200 microliter syringes, so that for the acetone we only fill them about half way.  It was noticeably easier to apply the isopropyl when the syringe only had 50 microliters.

* It may be helpful to have a strong, focused optical light to inspect the surface of the mirror.  Rana says that Garilynn might have such an optical fiber light that we could borrow.

ELOG V3.1.3-